40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7736     Entry time: Wed Nov 21 01:31:37 2012
Author: Koji, Ayaka 
Type: Update 
Category: Locking 
Subject: alignment on ETMX table 

Since the transmission beam on ETMXT camera seemed to be clipped, we checked the optics on ETMX table.

We aligned the lens so that it is orthogonal to the beam, then the beam shape looks fine.

output.nv12.bmp

Also we removed some an-used optics which were used for fiber input.

ELOG V3.1.3-